OpenStudy (anonymous):

solve the equation 5m/2m-1= 3/2

OpenStudy (anonymous):

5m/2m-1= 3/2 do you mean \[\frac{5m}{2m-1} = \frac{3}{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

2m-1 x 3 and 5m x 2?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

6m-3/10m

OpenStudy (anonymous):

this is the answer?

OpenStudy (anonymous):

6m-3 = 10m

OpenStudy (ryan123345):

now isolate the variable. subtract 6m on both sides

OpenStudy (anonymous):

ok I got it thanx guys